OpenStudy (anonymous):

Given that the legs of a right triangle measure a and b and the hypotenuse of the triangle measures c, what does the Pythagorean Theorem state about the sides of this triangle? a2 + c2 = b2 a2 − b2 = c2 2a + 2b = 2c a2 + b2 = c2

OpenStudy (korosh23):

The formula is a^2 + b^2= C^2

OpenStudy (korosh23):

 3^2 + 4^2=c^2 9+16= C^2 25= C^2 5= c

OpenStudy (korosh23):

This was an example. If my explanation was helpful, give me a medal and feel free to ask any question from me to clarify it much more. :)

OpenStudy (anonymous):

alright thank you i appreciate it